Calculate the molar mass of a liquid if 0.995g if it’s vapour occupy 250L at 200 degree cent and 45.0Torr


1
Expert's answer
2021-12-03T12:43:55-0500

"m=0.995 \; g \\

V = 250 \;L \\

T = 200 = 473 \; K \\

p=45.0 \;Torr = 0.0592 \; atm"

Ideal Gas Law

pV =nRT

R = 0.08206 L×atm/mol×K

"n = \frac{pV}{RT} \\

n = \frac{0.0592 \times 250}{0.08206 \times 473} = 0.381 \; mol \\

M = \frac{m}{n} \\

M = \frac{0.995}{0.381} = 2.61 \; g/mol"